# Quantum Phase Sensitivity with Generalized Coherent States Based on Deformed $su(1,1)$ and Heisenberg Algebras

### v1

@@ Abstract
We investigate the phase sensitivity of a Mach–Zehnder interferometer using a special class of generalized coherent states constructed from generalized Heisenberg and deformed $su(1,1)$ algebras. These states, derived from a perturbed harmonic oscillator with a four-parameter deformed spectrum, provide enhanced tunability and nonclassical features. The quantum Fisher information and its associated quantum Cramér-Rao bound are used to define the fundamental precision limits in phase estimation. We analyze the phase sensitivity under three realistic detection methods: difference intensity detection, single-mode intensity detection, and balanced homodyne detection. The performance of each method is compared with the quantum Cramér–Rao bound to evaluate their optimality. Our results demonstrate that, for suitable parameter regimes, these generalized coherent states enable phase sensitivities approaching the quantum limit, offering a flexible framework for precision quantum metrology.
@@ 1 Introduction
Interferometry is a technique based on the interference of waves. This technique is essential for precision measurement, quantum metrology and sensing, and is important for developing fundamental physics concepts [1]. Interferometric measurements of physical quantities like magnetic and gravitational fields fundamentally depend on detecting phase shifts in the interfering waves [2]. Due to their exceptional sensitivity to phase changes, interferometers are extensively utilised in precision measurement and metrology.
Quantum metrology is gaining widespread application across diverse fields, driven by advances in quantum technologies [9] quantum information theory [4, 5, 6, 7, 8]. As a fundamental aspect of quantum metrology, phase estimation has received considerable research attention. The sensitivity of optical interferometry is a research topic of interest and plays an essential role in various precision metrology applications [10], such as environmental sensing, gravitational wave detection [11, 12, 13], quantum technologies [14, 15], and gyroscopes [16].
@@ 2.1 Generalised Heisenberg algebra
and $\epsilon_{m+1}$ and $\epsilon_{m}$ denote two successive eigenvalues of Hamiltonian $\mathcal{H}$ such that $\epsilon_{m+1}=\varphi(\epsilon_{m})$ and $\epsilon_{0}$ represents the eigenvalue associated with the vacuum state. The operators $B$ and $B^{\dagger}$ are identified as the annihilation and creation operators of the GHA, respectively, with the vacuum condition $B|0\rangle=0$ being satisfied. Consequently, the eigenvalue $\varepsilon_{n}$ corresponds to the $n$-th iteration of $\varepsilon_{0}$ under the function $\varphi$, that is, $\varepsilon_{n}=\varphi^{n}(\varepsilon_{0})$. For the particular function $\varphi(\mathcal{H})=q\mathcal{H}+1$, the equations (1), and (2) recover the q-harmonic oscillator algebra. In the limit $q\rightarrow 1$, this algebra becomes the standard harmonic oscillator algebra. Moreover, the GHA has already been shown to be applicable to several physical systems characterized by a known spectrum, such as the Morse potential [50, 51], the square well potential [49], and the Poschl-Teller potential.
@@ 2.2 Generalised su(1,1) algebra
$\mathcal{H}|m\rangle=\epsilon_{m}|m\rangle,$ (12)
$G_{+}|m\rangle=\mathcal{N}_{m}|m+1\rangle,$ (13)
$G_{-}|m\rangle=\mathcal{N}_{m-1}|m-1\rangle,$ (14)
@@ 2.3 Construction of coherent states
In this subsection, we construct the coherent states corresponding to the GHA and the generalized su(1,1) algebra. We begin with the GHA coherent state, defined as the eigenstate of the annihilation operator
@@
The solution of the equation $\alpha_{m+1}N_{m}=\zeta\alpha_{m}$ for the coefficients of $|m\rangle$ is
@@ 2.4 Coherent State Construction under Harmonic Oscillator Perturbations
where $a$, $k$, $d$, and $e$ are real and non-zero constants satisfying the following conditions
@@
$|\zeta\rangle=e^{-\frac{|\zeta|^{2}}{2}}\sum_{m=0}^{\infty}\frac{\zeta^{n}}{\sqrt{n!}}|m\rangle.$ (38)
@@ 3 QFI and Phase Estimation with Generalized Coherent States in MZI
Figure 1: This is a schematic setup of a MZI, which includes two independent phase shifts and a beam splitter (BS1) with an adjustable transmission coefficient. The input state evolves through the interferometric paths before measurement.
We begin by examining the most general case in which phase shifts $\phi_{1}$ and $\phi_{2}$ are applied to the upper and lower arms of the interferometer, respectively. To circumvent the issue of accounting for unavailable resources, such as an external phase reference, the two-parameter estimation method is employed, as discussed in the literature [38, 39]. When an external phase reference is absent, the quantity of interest becomes the phase difference between the two interferometer arms, defined as $\phi_{d}=\phi_{1}-\phi_{2}$. Consequently, it is convenient to represent the quantum Fisher information matrix (QFIM) in terms of the parameters $\phi_{s/d}=\phi_{1}\pm\phi_{2}$. This 2×2 matrix [40, 41, 42] enables the joint estimation of $\phi_{s}$ and $\phi_{d}$ and is expressed as follows: representation
@@
$|\psi\rangle=e^{-i\frac{{}_{2}-\hat{m}_{3}}{2}\phi_{d}}e^{-i\frac{\hat{m}_{2}+\hat{m}_{3}}{2}\phi_{s}}|\psi_{23}\rangle,$ (41)
@@
The components of the QFIM, specifically $\mathcal{F}_{ss}$, $\mathcal{F}_{dd}$, and $\mathcal{F}_{sd}$—are obtained by applying the definition given in equation (40);
@@
Similar analytical expressions for the generalized SU(1,1) coherent state can be obtained by replacing the average photon number and its variance with the corresponding expressions computed for the state given in (LABEL:zsu).
@@ 4 Phase Sensitivity and Detection Schemes with Generalized Coherent State Inputs
We now complete the Mach-Zehnder interferometer by introducing the second beam splitter (BS2), defined by its transmission ($t^{\prime}$) and reflection ($r^{\prime}$) coefficients. Subsequently, we analyze the performance of three practical detection strategies: difference intensity detection, single-mode intensity detection, and balanced homodyne detection.
@@ 4.1 Phase Sensitivity under Intensity-Difference Detection
$\hat{b}_{4}$ $=e^{-i\phi_{2}}\left[(\cos{\tfrac{\kappa}{2}}\cos{\tfrac{\kappa^{\prime}}{2}}-\sin{\tfrac{\kappa}{2}}\sin{\tfrac{\kappa^{\prime}}{2}}e^{-i\phi})\,\hat{b}_{0}+i(\cos{\tfrac{\kappa}{2}}\sin{\tfrac{\kappa^{\prime}}{2}}e^{-i\phi}+\sin{\tfrac{\kappa}{2}}\cos{\tfrac{\kappa^{\prime}}{2}})\,\hat{b}_{1}\right],$ (83)
$\hat{b}_{5}$ $=e^{-i\phi_{2}}\left[i(\cos{\tfrac{\kappa}{2}}\sin{\tfrac{\kappa^{\prime}}{2}}+\sin{\tfrac{\kappa}{2}}\cos{\tfrac{\kappa^{\prime}}{2}}e^{-i\phi})\,\hat{b}_{0}+(\cos{\tfrac{\kappa}{2}}\cos{\tfrac{\kappa^{\prime}}{2}}e^{-i\phi}-\sin{\tfrac{\kappa}{2}}\sin{\tfrac{\kappa^{\prime}}{2}})\,\hat{b}_{1}\right].$ (84)
@@ 4.3 Phase Sensitivity via Balanced Homodyne Detection
$\Delta_{i}^{2}\hat{X}_{\phi_{L}}=$ $\frac{1}{4}-\frac{1}{2}\left(\cos^{2}{\frac{\kappa}{2}}\sin^{2}{\frac{\kappa^{\prime}}{2}}\cos 2\phi+\cos^{2}{\frac{\kappa^{\prime}}{2}}\sin^{2}{\frac{\kappa}{2}}+\frac{1}{2}\sin{\kappa}\sin{\kappa^{\prime}}\cos\phi\right)\left(|\langle\hat{b}^{2}_{1}\rangle_{i}|-|\langle\hat{b}_{1}\rangle_{i}|^{2}\right)$ (109)
$+\frac{1}{2}\left(\cos^{2}{\frac{\kappa}{2}}\sin^{2}{\frac{\kappa^{\prime}}{2}}+\cos^{2}{\frac{\kappa^{\prime}}{2}}\sin^{2}{\frac{\kappa}{2}}+\frac{1}{2}\sin{\kappa}\sin{\kappa^{\prime}}\cos\phi\right)\left(\langle\hat{m}_{1}\rangle_{i}-|\langle\hat{b}_{1}\rangle_{i}|^{2}\right),$ (110)
@@
$\Delta_{i}^{2}\hat{X}_{\phi_{L}}=$ $\frac{1}{4}-\frac{1}{2}\left((\cos^{2}{\frac{\kappa}{2}}\sin^{2}{\frac{\kappa^{\prime}}{2}}+\cos^{2}{\frac{\kappa^{\prime}}{2}}\sin^{2}{\frac{\kappa}{2}})\cos\phi+\frac{1}{2}\sin{\kappa}\sin{\kappa^{\prime}}\right)\left(|\langle\hat{b}^{2}_{1}\rangle_{i}|-|\langle\hat{b}_{1}\rangle_{i}|^{2}\right)$ (113)
$+\frac{1}{2}\left(\cos^{2}{\frac{\kappa}{2}}\sin^{2}{\frac{\kappa^{\prime}}{2}}+\cos^{2}{\frac{\kappa^{\prime}}{2}}\sin^{2}{\frac{\kappa}{2}}+\frac{1}{2}\sin{\kappa}\sin{\kappa^{\prime}}\cos\phi\right)\left(\langle\hat{m}_{1}\rangle_{i}-|\langle\hat{b}_{1}\rangle_{i}|^{2}\right),$ (114)
@@
$|\frac{\partial\langle\hat{X}_{\phi_{L}}\rangle}{\partial\phi}|=|\cos{\frac{\kappa^{\prime}}{2}}\sin{\frac{\kappa}{2}}-\cos{\frac{\kappa}{2}}\sin{\frac{\kappa^{\prime}}{2}}||\cos(\frac{\phi}{2})\langle\hat{b}_{1}\rangle_{i}|,$ (115)
@@
[…]
@@
Here, the interferometric configuration involves two arms: the signal arm traversing both beam splitters with intermediate phase shift $\phi_{1}$, and the local oscillator arm entering the balanced beam splitter of the homodyne setup.

### v2

@@ Abstract
We investigate the phase sensitivity of a Mach–Zehnder interferometer using a special class of generalized coherent states constructed from generalized Heisenberg and deformed $su(1,1)$ algebras. These states, derived from a perturbed harmonic oscillator with a four-parameter deformed spectrum, provide enhanced tunability and nonclassical features. The quantum Fisher information and its associated quantum Cramér-Rao bound are computed to define the fundamental precision limits in phase estimation. We analyze the phase sensitivity under three realistic detection methods: difference intensity detection, single-mode intensity detection, and balanced homodyne detection. The performance of each method is compared with the quantum Cramér–Rao bound to evaluate their optimality. Our results demonstrate that, for suitable parameter regimes, these generalized coherent states enable phase sensitivities approaching the quantum limit. This offers a flexible framework for precision quantum metrology and potential applications in quantum-enhanced sensing.
@@ 1 Introduction
Interferometry is a technique based on the interference of waves. This technique is essential for precision measurement, quantum metrology and sensing, and is important for developing fundamental physics concepts [1]. Interferometric measurements of physical quantities like magnetic and gravitational fields fundamentally depend on detecting phase shifts in the interfering waves [2]. Due to their exceptional sensitivity to phase changes, interferometers are extensively utilised in precision measurement and metrology. Specifically, quantum metrology is gaining widespread application across diverse fields, driven by advances in quantum technologies [3, 4] and quantum information theory [5, 6, 7, 8, 9]. As a fundamental aspect of quantum metrology, phase estimation has received considerable research attention. The sensitivity of optical interferometry is a research topic of interest and plays an essential role in various precision metrology applications [10, 11, 12, 13, 14], such as environmental sensing, gravitational wave detection [15, 16, 17], quantum technologies [18, 19], and gyroscopes [20].
@@ 2.1 Generalised Heisenberg algebra
and $\epsilon_{m+1}$ and $\epsilon_{m}$ denote two successive eigenvalues of Hamiltonian $\mathcal{H}$ such that $\epsilon_{m+1}=\varphi(\epsilon_{m})$ and $\epsilon_{0}$ represents the eigenvalue associated with the vacuum state. The operators $B$ and $B^{\dagger}$ are identified as the annihilation and creation operators of the GHA, respectively, with the vacuum condition $B|0\rangle=0$ being satisfied. Consequently, the eigenvalue $\varepsilon_{n}$ corresponds to the $n$-th iteration of $\varepsilon_{0}$ under the function $\varphi$, that is, $\varepsilon_{n}=\varphi^{n}(\varepsilon_{0})$. For the particular function $\varphi(\mathcal{H})=q\mathcal{H}+1$, the equations (1), and (2) recover the q-harmonic oscillator algebra. In the limit $q\rightarrow 1$, this algebra becomes the standard harmonic oscillator algebra. Moreover, the GHA has already been shown to be applicable to several physical systems characterized by a known spectrum, such as the Morse potential [43], the square well potential [44], and the Poschl-Teller potential [45].
@@ 2.2 Generalised su(1,1) algebra
$\mathcal{H}\left|m\right\rangle=\epsilon_{m}\left|m\right\rangle,$ (12)
$G_{+}\left|m\right\rangle=\mathcal{N}_{m}\left|m+1\right\rangle,$ (13)
$G_{-}\left|m\right\rangle=\mathcal{N}_{m-1}\left|m-1\right\rangle,$ (14)
@@ 2.3 Construction of coherent states
In fact, coherent states are seen as the closest quantum analogue to classical radiation fields, reproducing many of the wave properties of light [57]. Introduced by Glauber [46], they have acquired a central place in quantum optics, particularly after Glauber showed them to be the eigenstates of the harmonic oscillator’s annihilation operator while simultaneously realizing the minimal bound of the Heisenberg uncertainty relation [58, 59, 60]. Subsequently, Perelomov generalized Glauber’s approach to the framework of arbitrary Lie groups [41, 61]. In this formalism, a generalized coherent state is obtained simply by applying a generalized displacement operator defined as the unitary exponentiation of the considered Lie algebra to an extremal state. In this subsection, we construct the coherent states corresponding to the GHA and the generalized su(1,1) algebra. We begin with the GHA coherent state, defined as the eigenstate of the annihilation operator
@@
The solution of the equation $\alpha_{m+1}N_{m}=\zeta\alpha_{m}$ for the coefficients of $\left|m\right\rangle$ is
@@
The deformation parameters encoded in $N_{m}$ or $\mathcal{N}_{m}$ modify the oscillator’s energy-level spacing, introducing nonlinearities into the system. This allows interpolation between standard Glauber coherent states describing classical like light fields and deformed states with enhanced nonclassical properties such as squeezing and sub-Poissonian statistics. Glauber originally introduced coherent states to represent fields with Poissonian photon number statistics and minimal uncertainty [46]. Similar nonclassical features have been studied in hypergeometric states [47]. In the limit where deformation vanishes, our formalism reduces to the standard harmonic oscillator coherent states, preserving the connection to quantum optics.
@@ 2.4 Coherent State Construction under Harmonic Oscillator Perturbations
where $a$, $k$, $d$, and $e$ are real and non-zero constants satisfying the following conditions [48]
@@
$|\zeta\rangle=\exp\left[{-\frac{|\zeta|^{2}}{2}}\right]\sum_{m=0}^{\infty}\frac{\zeta^{n}}{\sqrt{n!}}|m\rangle.$ (38)
The parameters $a$, $d$, and $e$ determine how strongly the oscillator departs from the harmonic potential, thereby controlling squeezing, photon-number fluctuations, and the QFI relevant to phase sensitivity. Larger nonlinearities enhance quantum features useful for metrology, while the ordinary oscillator limit recovers coherent states with Poissonian statistics [48, 49].
@@ 3 QFI and Phase Estimation with Generalized Coherent States in MZI
Figure 1: This is a schematic setup of a MZI, which includes two independent phase shifts and two beam splitters (BS1 and BS2) with an adjustable transmission coefficient. The input state evolves through the interferometric paths before measurement.
We begin by examining the most general case in which phase shifts $\phi_{1}$ and $\phi_{2}$ are applied to the upper and lower arms of the interferometer, respectively. To circumvent the issue of accounting for unavailable resources, such as an external phase reference, the two-parameter estimation method is employed, as discussed in the literature [50, 51]. When an external phase reference is absent, the quantity of interest becomes the phase difference between the two interferometer arms, defined as $\phi_{d}=\phi_{1}-\phi_{2}$. Consequently, it is convenient to represent the quantum Fisher information matrix (QFIM) in terms of the parameters $\phi_{s/d}=\phi_{1}\pm\phi_{2}$. This 2×2 matrix [52, 53, 54] enables the joint estimation of $\phi_{s}$ and $\phi_{d}$ and is expressed as follows:
@@
$|\psi\rangle=\exp\left[{-i\frac{{}_{2}-\hat{m}_{3}}{2}\phi_{d}}\right]\exp\left[{-i\frac{\hat{m}_{2}+\hat{m}_{3}}{2}\phi_{s}}\right]|\psi_{23}\rangle,$ (41)
@@
The components of the QFIM, specifically $\mathcal{F}_{ss}$, $\mathcal{F}_{dd}$, and $\mathcal{F}_{sd}$ are obtained by applying the definition given in equation (40);
@@
Similar analytical expressions for the generalized SU(1,1) coherent state can be obtained by replacing the average photon number and its variance with the corresponding expressions computed for the state given in Eq. (22). Based on these derivations, the corresponding QFIs and QCRBs for the different phase estimation scenarios are summarized in Table 1.
Scenario QFI QCRB
Two-parameter estimation (a) $F^{(a)}_{i}=\sin^{2}\kappa\,\langle\hat{m}_{1}\rangle_{i}$ $\Delta\varphi^{(a)}_{\text{QCRB}}=\dfrac{1}{\sqrt{F^{(a)}_{i}}}$
Single-arm phase shift (b) $F^{(b)}_{i}=4\cos^{4}\!\dfrac{\kappa}{2}\,\Delta^{2}\hat{m}_{1}+\sin^{2}\kappa\,\langle\hat{m}_{1}\rangle_{i}$ $\Delta\varphi^{(b)}_{\text{QCRB}}=\dfrac{1}{\sqrt{F^{(b)}_{i}}}$
Symmetric phase shift (c) $F^{(c)}_{i}=\Big(\cos^{4}\!\dfrac{\kappa}{2}+\sin^{4}\!\dfrac{\kappa}{2}\Big)\Delta^{2}\hat{m}_{1}+\tfrac{1}{2}\sin^{2}\kappa\,\langle\hat{m}_{1}\rangle_{i}$ $\Delta\varphi^{(c)}_{\text{QCRB}}=\dfrac{1}{\sqrt{F^{(c)}_{i}}}$
Table 1: Summary of the analytical results for the QFI and the associated QCRB for different phase estimation scenarios.
@@ 4 Phase Sensitivity and Detection Schemes with Generalized Coherent State Inputs
We now complete the Mach-Zehnder interferometer by introducing the second beam splitter (BS2), as depicted in Figure 2, defined by its transmission ($t^{\prime}$) and reflection ($r^{\prime}$) coefficients. Subsequently, we analyze the performance of three practical detection strategies: difference intensity detection, single-mode intensity detection, and balanced homodyne detection.
@@ 4.1 Phase Sensitivity under Intensity-Difference Detection
$\hat{b}_{4}$ $=\exp\left[{-i\phi_{2}}\right]\left[(\cos{\tfrac{\kappa}{2}}\cos{\tfrac{\kappa^{\prime}}{2}}-\sin{\tfrac{\kappa}{2}}\sin{\tfrac{\kappa^{\prime}}{2}}e^{-i\phi})\,\hat{b}_{0}+i(\cos{\tfrac{\kappa}{2}}\sin{\tfrac{\kappa^{\prime}}{2}}e^{-i\phi}+\sin{\tfrac{\kappa}{2}}\cos{\tfrac{\kappa^{\prime}}{2}})\,\hat{b}_{1}\right],$ (83)
$\hat{b}_{5}$ $=\exp\left[{-i\phi_{2}}\right]\left[i(\cos{\tfrac{\kappa}{2}}\sin{\tfrac{\kappa^{\prime}}{2}}+\sin{\tfrac{\kappa}{2}}\cos{\tfrac{\kappa^{\prime}}{2}}e^{-i\phi})\,\hat{b}_{0}+(\cos{\tfrac{\kappa}{2}}\cos{\tfrac{\kappa^{\prime}}{2}}e^{-i\phi}-\sin{\tfrac{\kappa}{2}}\sin{\tfrac{\kappa^{\prime}}{2}})\,\hat{b}_{1}\right].$ (84)
@@ 4.3 Phase Sensitivity via Balanced Homodyne Detection
$\Delta_{i}^{2}\hat{X}_{\phi_{L}}=$ $\frac{1}{4}-\frac{1}{2}\left(\cos^{2}{\frac{\kappa}{2}}\sin^{2}{\frac{\kappa^{\prime}}{2}}\cos 2\phi+\cos^{2}{\frac{\kappa^{\prime}}{2}}\sin^{2}{\frac{\kappa}{2}}+\frac{1}{2}\sin{\kappa}\sin{\kappa^{\prime}}\cos\phi\right)\left(|\langle\hat{b}^{2}_{1}\rangle_{i}|-|\langle\hat{b}_{1}\rangle_{i}|^{2}\right)$
$+\frac{1}{2}\left(\cos^{2}{\frac{\kappa}{2}}\sin^{2}{\frac{\kappa^{\prime}}{2}}+\cos^{2}{\frac{\kappa^{\prime}}{2}}\sin^{2}{\frac{\kappa}{2}}+\frac{1}{2}\sin{\kappa}\sin{\kappa^{\prime}}\cos\phi\right)\left(\langle\hat{m}_{1}\rangle_{i}-|\langle\hat{b}_{1}\rangle_{i}|^{2}\right).$ (109)
@@
$\Delta_{i}^{2}\hat{X}_{\phi_{L}}=$ $\frac{1}{4}-\frac{1}{2}\left((\cos^{2}{\frac{\kappa}{2}}\sin^{2}{\frac{\kappa^{\prime}}{2}}+\cos^{2}{\frac{\kappa^{\prime}}{2}}\sin^{2}{\frac{\kappa}{2}})\cos\phi+\frac{1}{2}\sin{\kappa}\sin{\kappa^{\prime}}\right)\left(|\langle\hat{b}^{2}_{1}\rangle_{i}|-|\langle\hat{b}_{1}\rangle_{i}|^{2}\right)$
$+\frac{1}{2}\left(\cos^{2}{\frac{\kappa}{2}}\sin^{2}{\frac{\kappa^{\prime}}{2}}+\cos^{2}{\frac{\kappa^{\prime}}{2}}\sin^{2}{\frac{\kappa}{2}}+\frac{1}{2}\sin{\kappa}\sin{\kappa^{\prime}}\cos\phi\right)\left(\langle\hat{m}_{1}\rangle_{i}-|\langle\hat{b}_{1}\rangle_{i}|^{2}\right).$ (112)
@@
$|\frac{\partial\langle\hat{X}_{\phi_{L}}\rangle}{\partial\phi}|=|\cos{\frac{\kappa^{\prime}}{2}}\sin{\frac{\kappa}{2}}-\cos{\frac{\kappa}{2}}\sin{\frac{\kappa^{\prime}}{2}}||\cos(\frac{\phi}{2})\langle\hat{b}_{1}\rangle_{i}|.$ (113)
@@
The analytical results for the phase sensitivities obtained with the three considered detection schemes are summarized in Table 2.
Detection Scheme Phase Sensitivity
Intensity-difference detection $\Delta\varphi^{\text{df}}_{i}=\dfrac{\Delta\hat{N}_{d}}{\sin\kappa\sin\kappa^{\prime}|\sin\varphi|\langle\hat{m}_{1}\rangle_{i}}$
Single-mode intensity detection $\Delta\varphi^{\text{sing}}_{i}=\dfrac{2\Delta\hat{m}_{4}}{\sin\kappa\sin\kappa^{\prime}|\sin\varphi|\langle\hat{m}_{1}\rangle_{i}}$
Balanced homodyne detection (b) $\Delta\varphi^{(b)}_{\text{hom},i}=\dfrac{\sqrt{\Delta^{2}\hat{X}^{\varphi_{L}}_{i}}}{\cos\!\dfrac{\kappa^{\prime}}{2}\,\sin\!\dfrac{\kappa}{2}\,|\cos(\varphi)\langle\hat{b}_{1}\rangle_{i}|}$
Balanced homodyne detection (c) $\Delta\varphi^{(c)}_{\text{hom},i}=\dfrac{\sqrt{\Delta^{2}\hat{X}^{\varphi_{L}}_{i}}}{\big|\cos\!\dfrac{\kappa^{\prime}}{2}\,\sin\!\dfrac{\kappa}{2}-\cos\!\dfrac{\kappa}{2}\,\sin\!\dfrac{\kappa^{\prime}}{2}\big|\,|\cos(\tfrac{\varphi}{2})\langle\hat{b}_{1}\rangle_{i}|}$
Table 2: Summary of the analytical results for the phase sensitivities obtained with different detection schemes.
@@
Here, the interferometric configuration involves two arms: the signal arm traversing both beam splitters with intermediate phase shift $\phi_{1}$, and the local oscillator arm entering the balanced beam splitter of the homodyne setup. Previous works on q-deformed and non-linear coherent states have shown that deformation enhances squeezing and sub-Poissonian statistics, enabling phase sensitivities beyond the SNL [62, 63]. Our contribution goes further by employing generalized Heisenberg and generalized su(1,1) algebras with multiple deformation parameters, offering greater tunability and control of non-classical properties. This richer algebraic structure optimizes quantum Fisher information and phase sensitivity, thereby extending and complementing previous deformed-algebra approaches in quantum metrology.